Circular No. 8624

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2005 U1 (READ) M. T. Read reports the discovery of a diffuse comet with a tail ∼ 20′′

long toward the southwest on CCD mosaic images taken with the Spacewatch 0.9-m f/3 reflector on Oct. 24.3 UT (discovery observation tabulated below). The cometary appearance has been confirmed on Spacewatch images taken with the 1.8-m f/2.7 reflector by T. H. Bressi on Oct. 25.35–25.36 (20′′ tail extending westward), by T. Gehrels on Oct. 26.30–26.31 (6′′ coma and tail at least 10′′ long in p.a. 260^{o}) and 27.28–27.29 (small, elongated coma and tail > long in p.a. 260^{o}), and by J. V. Scotti on Oct. 28.4 (slightly diffuse, with 0′.37 tail in p.a. 249^{o}). E. J. Christensen reports that images taken with the Mt. Lemmon 1.5-m reflector on Oct. 28.3 show a 6′′ coma and a 10′′ tail in p.a. 240^{o}–250^{o}.

2005 UT  α_{2000}  δ_{2000}  Mag.
Oct. 24.32776  3^{h}21^{m}11.^{s}94  +18^{o}09′17′′.8  20.2

The available astrometry (including prediscovery Catalina observations on Oct. 23), the following preliminary orbital elements, and an ephemeris appear on MPEC 2005-U74.

T = 2005 July 7.146 TT  ω = 316.^{o}571
e = 0.31063  Ω = 51.812 }2000.0
q = 2.25603 AU  i = 1.188
a = 3.27260 AU n^{o} = 0.166481 P = 5.92 years

COMET C/2004 U1 (LINEAR) J. Young, Table Mountain Observatory, reports the discovery of a very

faint second component (of total mag ∼ 20.0), moving at the same rate as the primary component of C/2004 U1 (total mag ∼ 18.5), on stacked CCD images taken by D. Mayes with the 0.6-m f/16 Cassegrain reflector on Oct. 20.3 and 21.3 UT. The companion appears to be a well-defined spot along what may be a faint tail of the primary. The astrometry has been published on MPEC 2005-U66.

COMET C/2005 T4 (SWAN) Magnitude estimates and coma diameters from sets of three co-added

3-min exposures by T. Lovejoy (Thornlands, Qld., 100-mm f/2.8 telephoto lens): Oct. 9.378 UT, 10.5, 2′–3′; 18.372, 12, –.

2005 October 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green